There is a bag filled with 3 blue, 2 red and 4 green marbles. A marble is taken at random from the bag, the colour is noted and then it is not replaced. Another marble is taken at random. What is the probability of getting 2 different colours?


1
Expert's answer
2022-05-03T16:51:10-0400

We consider two possibilities:

  1. The order of marbles does not matter. I.e., it does not matter, which marble was taken first and which marble was taken second. There are: "C_9^2=\frac{9\cdot8}2=36" different ways to choose "2" marbles from "9". Denote by the capital letter a color of the marble. The following combinations correspond to two different colors: "BR, RB,BG,GB, RG,GR."We get: "\frac{6+6}2=6" ways to choose a blue and a red marble, "\frac{12+12}{2}=12" ways to choose a blue and a green marble and "\frac{8+8}{2}=8" ways to choose a red and a green marble. A division by "2" is present, since the order of marbles does not matter. The probability to get two marbles with different colors is: "\frac{6+12+8}{36}=\frac{26}{36}=\frac{13}{18}"
  2. The order of marbles matters. There are "9\cdot8=72" ways to choose "2" marbles from "9". The following combinations correspond to two different colors: "BR, RB,BG,GB, RG,GR". Find the respective probabilities for these combinations: "P(BR)=P(RB)=\frac{6}{72}", "P(BG)=P(GB)=\frac{12}{72}", "P(RG)=P(GR)=\frac{8}{72}". The probability is: "p=\frac{12+24+16}{72}=\frac{52}{72}=\frac{13}{18}"

The answer is: "\frac{13}{18}."